OpenStudy (anonymous):

A Farmer has 142 animals chickens and goats if the animals have a total of 496 legs what ise number of chickens?

OpenStudy (anonymous):

let chicken be x goat be y total animals are x+y=142

OpenStudy (anonymous):

legs of chicken is =2x legs of goat =4y 2x+4y=496

OpenStudy (anonymous):

Let chickens be x and goat be y. As we know that chicken have two legs and goat have four legs. 2x + 4y = 496. Total animals = 142 i.e. x+y = 142

OpenStudy (anonymous):

solve the 2 equations....:)

OpenStudy (aripotta):

system of equationsssss whooooooooop

OpenStudy (anonymous):

x + y = 142 x = 142 - y ----- (1) 2x + 4y = 196 ---- (2) Substituting equation (1) on (2) 2(142-y) + 4y = 192 by solving these you will get the value of y then put the value in eqn. (1), you will get the x.

OpenStudy (anonymous):

x=36

OpenStudy (aripotta):

was typing whole answer, then saw above.........

OpenStudy (aripotta):

life = wasted.

OpenStudy (anonymous):

it ok sir be happy:)

OpenStudy (aripotta):

me gusta system of equations.